OpenStudy (anonymous):

PLEASE HELP! 7v-1/2=6/5v-3/5

OpenStudy (anonymous):

5(7v-1/2) = 5(6/5v-3/5) 35v-5/2=6v-3 29v= 3+5/2 v= 11/2/29 v=11/58

OpenStudy (anonymous):

that is incorrect :/

myininaya (myininaya):

I like to get rid of fractions so multiply both sides by 10 so we have 70v-5=12v-6 70v-12v=-6+5 58v=-1 v=-1/58

OpenStudy (anonymous):

Yep, it should be -1/58

OpenStudy (anonymous):

ok what about 1/5w+4/3=7/3w-4 ??

myininaya (myininaya):

like i said i like to get rid of the fractions multiply both sides by 15 3w+20=35w-60 3w-35w=-60-20 -32w=-80 w=80/32=5/2

OpenStudy (anonymous):

what do u mean by both sides?

OpenStudy (anonymous):

what do u multiply by 15?

OpenStudy (anonymous):

By both sides myininaya probably means to multiply the terms on the right and left of the = sign by that amount.

OpenStudy (anonymous):

i'm confused :/

OpenStudy (anonymous):

can you show me with this example? 5/4*u+3/4=1/3*u+7

myininaya (myininaya):

you don't have to get rid of the fractions you could just work with them w/5-7w/3=-4-4/3 (3w-35w)/15=(-12-4)/3 -32w/15=-16/3 w=(-16/3)(15/-32)=5/2

OpenStudy (anonymous):

see like this one....u multiply every single fraction by 15 to get rid of the fractions?

myininaya (myininaya):

do u see that we have fractions in this new problem right?

OpenStudy (anonymous):

yess

myininaya (myininaya):

what do you see on the bottoms of these fractions? you see the numbers 4 and 3, right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

4*3=12 so multiply both sides by 12

myininaya (myininaya):

what do you get?

OpenStudy (anonymous):

what number do i multiply 12 with?

myininaya (myininaya):

what is 12*5u/4 what is 12*(3/4) what is 12*(u/3) what is 12*7

OpenStudy (anonymous):

ohhh okayy

myininaya (myininaya):

15u 9 4u 84

myininaya (myininaya):

so we have 15u+9=4u+84

OpenStudy (anonymous):

k i think i can do it from here. let me see if i can figure it out

myininaya (myininaya):

k brb

OpenStudy (anonymous):

is it 6.81 ???

OpenStudy (anonymous):

6 81/100 ??? i dont think that's right :/

OpenStudy (anonymous):

or 17/11 ??

myininaya (myininaya):

you still there?

myininaya (myininaya):

15u-4u=84-9 11u=75 u=75/11